# Non-transitive pseudo-Anosov flows

### v1

@@ 2.1. Structure of the bifoliated planes
Here is where a little extra care is needed compared to the case of transitive Anosov-like actions. The assertion above is true for any $l^{\prime\prime}$ in $D(l)$ fixed by some non trivial element $h\in G$. So, thanks to Axiom (A2), we can find such a leaf $l^{\prime\prime}$ that intersects $\mathcal{F}^{+}(x)$. Calling $x^{\prime\prime}$ the fixed point of $h$ on $l^{\prime\prime}$, we will first show that $\mathcal{F}^{+}(x^{\prime\prime})$ intersects $l^{\prime}$ [^3] This is the only argument that needs to be added compared to [Bar95a] or [BFM22], since in the transitive case, one can choose $x^{\prime\prime}$ close to $x$ so that property is automatically satisfied.. Consider the set $S$ of $\mathcal{F}^{+}$ leaves intersecting both $L^{\prime\prime}$ and $l^{\prime}$. If $\mathcal{F}^{+}(x^{\prime\prime})$ is not in this set, then it means that there is a leaf $l_{1}^{+}$ in $\partial S$ that separates $\mathcal{F}^{+}(x^{\prime\prime})$ from $S$. Since $\alpha(l^{\prime})$ and $\omega(l^{\prime})$ are not in $S$ either, there are also leaves $l_{\alpha}^{+},l_{\omega}^{+}\subset\partial S$ that separates $S$ from $\alpha(l^{\prime})$ and $\omega(l^{\prime})$ respectively (we may a priori have $\alpha(l^{\prime})\subset\partial S$ in which case $\alpha(l^{\prime})=l_{\alpha}^{+}$ and same for $\omega(l^{\prime})$). Then $l_{1}^{+}$ is not separated with either $l_{\alpha}^{+}$ or $l_{\omega}^{+}$, contradicting the fact that $\mathcal{L}^{+}$ is homeomorphic to $\mathbb{R}$. Hence, $\mathcal{F}^{+}(x^{\prime\prime})$ intersects $l^{\prime}$ and one finishes the proof as in [Bar95a] or [BFM22]. ∎
@@
This was proved for flows by Barbot [Bar95a, Théorème 2.7], and for transitive Anosov-like actions in [BFM22, Proposition 2.7], the proof there does not use the full strength of transitivity and carries over in our setting without modification. In the case of flows, this is equivalent to the fact that any (pseudo)-Anosov flow with trivial orbit space is the suspension flow of a hyperbolic linear map of the torus.
@@ 2.2. Axiom (A4) and non-separated leaves
Axiom (A4) states that any non-separated leaf is fixed by some nontrivial element of $G$. Our next lemma improves this to show that for any maximal set of pairwise non-separated leaves, there exists a single (nontrivial) element $g\in G$ fixing all the leaves in that set. In particular this shows that Axiom (A4) implies the (a priori stronger) corresponding axiom in [BFM22].
@@
Take $l_{0}\in S$ and $g\in G$ fixing $l_{0}$. If $g$ is trivial there is nothing to show; so we assume $g$ nontrivial. By Axiom (A1), there exists a unique $x_{0}\in l_{0}$ fixed by $g$. Up to taking a power, we further assume that that $g$ fixes all the rays of ±$\mathcal{F}^{\pm}(x_{0})\smallsetminus\{x_{0}\}$.
@@
Pick $l_{0}^{\prime}\in S$ another leaf. By Observation 2.11, there exists $f_{0}\in\mathcal{F}^{-}$ that makes a perfect fit with $l_{0}$ and separates $l_{0}$ from $l^{\prime}_{0}$. As $g$ fixes every ray of $x_{0}$, it also fixes $f_{0}$ and Axiom (A1) again gives the existence of a fixed point $y_{0}\in f_{0}$. Let $r_{0}$ be the ray of $\mathcal{F}^{+}(y_{0})\smallsetminus\{y_{0}\}$ that is contained in the side of $f_{0}$ that contains $l^{\prime}_{0}$. Call $S_{0}$ the subset of $S$ on that same side of $f_{0}$. So in particular, $l^{\prime}_{0}\in S_{0}$.
@@
Let $z$ be as above. There exists a unique leaf $f_{1}\in\mathcal{F}^{-}$ such that it, or one of its face in the singular case, is in $\partial(P\smallsetminus\mathcal{F}^{-}(r_{0}))$ and separates $z$ from $x_{0}$. We claim that $f_{1}\cap r_{0}=\emptyset$. This is automatic if $f_{1}$ is non-singular, but assume not and call $t_{0}=f_{1}\cap r_{0}$. Then, by construction, for all $t>t_{0}$ in $r_{0}$ (where $>$ designate the positive direction of the ray), we have $\mathcal{F}^{-}(t)\cap S_{0}=\emptyset$. However, since $S$ consists of pairwise non-separated leaves, there exists $l^{+}$ that intersects both $f_{0}$ and $\mathcal{F}^{-}(z)$. But (up to replacing $g$ by $g^{-1}$) the sequence $g^{n}l^{+}$ contains $\mathcal{F}^{+}(y_{0})$ in its limit as $n\to+\infty$. In particular, we would get $t>t_{0}$ in $r_{0}$ such that $\mathcal{F}^{-}(t)\cap S_{0}\neq\emptyset$, a contradiction.
@@ 2.3. Product and scalloped regions
Suppose $U$ is an infinite product region in $P$, bounded by an interval $I$. Up to switching the labels $+$ and --, we assume $I$ lies in a leaf of $\mathcal{F}^{+}$. By axiom (A2) there is a leaf $l^{-}$ of $\mathcal{F}^{-}$ intersecting $U$ and fixed by some $g\in G$. Let $x$ be the unique fixed point of $g$ on $l^{-}$. Then $\bigcup_{k\in\mathbb{Z}}g^{k}(U)$ contains an infinite product region bounded by $\mathcal{F}^{+}(x)$. Let $y_{i}$ be points along $\mathcal{F}^{+}(x)$ such that the sequence $y_{i}$ eventually leaves every compact set, and for each $i$ let $r_{i}$ be a ray of $\mathcal{F}^{-}(y_{i})$ contained in $U$.
@@
Lemma 2.17 (Lemma 2.23 and Prop. 2.24 in [BFM22], see also [Fen95]).
@@
The proof of Lemma 2.23 and Prop. 2.24 in [BFM22] applies directly in this setting, given Proposition 2.16 and Proposition 2.10. ∎
We introduce another definition (from [BFM22]).
@@
It turns out that infinite lines of lozenges are special: They are part of what is called a scalloped region. A scalloped region is a trivially foliated open set $U\subset P$ whose boundary $\partial U$ consists in four infinite families of pairwise non-separated leaves. In particular, a scalloped region can be realized as two distinct bi-infinite lines of lozenges: one sharing sides along leaves of $\mathcal{F}^{+}$ and the other sharing $\mathcal{F}^{-}$-sides. See Figure 2. A formal definition can be found in e.g., [BFM22, Definition 2.31].
@@
The fact that infinite lines are contained in bi-infinite lines is given by Lemma 2.32 of [BFM22], the second assertion is given by Lemma 2.33 of [BFM22]. Note that neither of these proofs use the transitivity axioms, hence apply directly in our more general setting. ∎
@@ 2.4. Finite branching
$\mathrm{Pivot}:=\{p\in P\mid\text{a half leaf of }p\text{ is the side of two lozenges}\}$
@@
By definition the set of singular points is discrete and it is also closed since any nonsingular point has a product foliated neighborhood.
Suppose $x\in\mathrm{Pivot}$. Since $\mathrm{Sing}$ is closed and discrete, we can assume $x$ is nonsingular. We will find a neighborhood of $x$ containing no other pivot points. First, we claim that any point inside the lozenges sharing a side along ±$\mathcal{F}^{\pm}(x)$ (as in the definition of pivot) cannot be pivot points. To see this, consider such a point $y$. By the “non-corner criterion” [BFM22, Lemma 2.29], only one quadrant of $y$ can contain a lozenge with $y$ as its corner, namely the quadrant opposite to the one containing $x$. Thus, $y$ is not a pivot, as desired. Now, suppose we have some other pivot point $z_{1}$ in a product-foliated neighborhood containing $x$ (otherwise, we are already done). Up to switching local transverse orientations in that neighborhood, and reversing the names of $\mathcal{F}^{\pm}$, we can assume that $x$ is the corner of two adjacent lozenges in the half-space above $\mathcal{F}^{+}(x)$, and that $z_{1}$ is in the lower right quadrant of $x$. If $z_{1}$ is a pivot, there are a priori two possible configurations for the lozenges of $z_{1}$, see Figure 3. In both configurations, $z_{1}$ has a lozenge in its lower right quadrant. By the non-corner criterion, no point in the rectangle bounded by ±$\mathcal{F}^{\pm}(z_{1})$, $\mathcal{F}^{-}(x)$ and any $\mathcal{F}^{+}$-leaf slightly above $\mathcal{F}^{+}(x)$, can be a pivot point. Thus, any sequence of pivot points approaching a point $z$ would contain another pivot point $z_{2}$ which must be in the lower left quadrant of $x$ and $z_{2}$ admits a lozenge in its lower left quadrant.
[…]
Figure 3. Configurations for pairs of pivot lozenges at $x$ and $z_{1}$
We conclude that the rectangle bounded by ±$\mathcal{F}^{\pm}(z_{1})$, $\mathcal{F}^{-}(z_{2})$ and any $\mathcal{F}^{+}$-leaf slightly above $\mathcal{F}^{+}(x)$ cannot contain any pivot aside from $x$. Hence we proved that any point has a trivially foliated neighborhood containing at most one pivot. This proves that $\mathrm{Pivot}$ is closed and discrete. ∎
@@ 3. Smale classes and topological transitivity
In this section, we generalize the Smale order and basic sets for Anosov flows to the context of an Anosov-like action of a group $G$ on bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$. In particular, we show how the classical Smale order and basic sets (see [Sma67]) can be explicitly detected in the orbit space, that Smale classes have a product structure, and we prove Theorems 1.7 and 1.8. Later, in Section 4, we will also recover the analogue of tori separating basic sets from the orbit space.
@@
The regular set $\mathcal{R}_{G}$ is the subset of $\overline{\mathrm{Fix}}_{G}$ consisting of points $x\in\overline{\mathrm{Fix}}_{G}$ such that no leaf of ±$\mathcal{F}^{\pm}(x)$ is a singular leaf. Equivalently, $\mathcal{R}_{G}$ is obtained from $\overline{\mathrm{Fix}}_{G}$ by removing all prongs and their leaves.
@@
However, even for actions coming from pseudo-Anosov flows on compact manifolds, there may be differences between chain-recurrent and non-wandering sets. In this case $\overline{\mathrm{Fix}}_{G}$ corresponds to the projection of the non-wandering set. The non-wandering set can be strictly smaller than the chain recurrent set, as shown by the examples given in Propositions 8.8, and 8.9.
@@
In the case of the orbit space of an Anosov flow, the relation $\leq_{G}$ is defined and a transitive relation on all of $\overline{\mathrm{Fix}}_{G}$, giving the classical Smale order, and the Smale classes are basic sets ([Sma67], or see e.g., [FH19, Section 5.3]). Transitivity of the relation $\leq_{G}$ on $\overline{\mathrm{Fix}}_{G}$ also holds whenever $\mathcal{F}^{\pm}$ are true foliations, without singularities, by Lemma 3.4 below. However, in the presence of prongs, the property of having nonempty intersection is not open in the space of pairs of leaves in $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$ and transitivity may fail for points in $\overline{\mathrm{Fix}}_{G}$ – see Proposition 8.7. This is why we introduced the regular set $\mathcal{R}_{G}$; it is the set where $\leq_{G}$ always induces a partial order on Smale classes, as we show in the next lemma.
@@
The relation $\leq_{G}$ is reflexive and transitive on $\mathcal{R}_{G}$. Consequently, $\leq_{G}$ induces a partial order on Smale classes in $\mathcal{R}_{G}$.
@@
Definition 3.5 (Extremal classes, bounded actions).
We call a Smale class extremal if it is either minimal or maximal for the Smale order. We say that an Anosov-like action $G$ is Smale-bounded if for any Smale class $\Lambda_{i}$, there exist extremal Smale classes $\Lambda_{j},\Lambda_{k}$ such that $\Lambda_{j}\leq_{G}\Lambda_{i}\leq_{G}\Lambda_{k}$.
Obviously, any action with only finitely many Smale classes (such as those coming from pseudo-Anosov flows on compact 3-manifolds) is Smale-bounded. As a special case we have:
Proposition 3.6.
If the action of $G$ on $P$ is co-compact, then $(P,\leq_{G})$ has finitely many Smale classes, hence is Smale-bounded.
Proof.
Let $\Lambda_{i}$ be a Smale class. Each $x\in\Lambda_{i}$ has a neighborhood $U_{x}$ that is product foliated, so any $y\in U_{x}$ satisfies $y\sim_{G}x$. Suppose $G$ acts cocompactly on $P$. Since Smale classes are $G$-invariant, each is represented in a compact fundamental domain. By the previous remark, there can be no accumulation points of distinct classes, so there are only finitely many. ∎
There exist example of Anosov like actions that are non-cocompact but are Smale bounded, see Example 8.2. Later, starting in Section 6, we will restrict our attention to Smale-bounded actions. However, it is also easy to build examples of actions that are not bounded, even in the flows setting, if one allows for non-compactness, see Example 8.4
Our next goal is to give an equivalent characterization of Smale classes that is easier to work with. For this, we use the following definition and Lemma from [BFM22]
@@
$\mathcal{F}^{+}(a)\cap\mathcal{F}^{-}(b)\neq\emptyset\text{ and }\mathcal{F}^{-}(a)\cap\mathcal{F}^{+}(b)\neq\emptyset.$
Our next goal is to show that, as in the classical case of hyperbolic basic sets, the Smale classes of an Anosov-like action have a product structure, in the same sense as hyperbolic sets – see e.g. [FH19, Definition 6.2.6].
@@
Let $a,b\in\overline{\mathrm{Fix}}_{G}$ satisfy $x=\mathcal{F}^{+}(a)\cap\mathcal{F}^{-}(b)\neq\emptyset$. Let $r_{a}$ and $r_{b}$ be the rays of $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{-}(b)$ (respectively) containing $x$, and let $f_{a},f_{b}$ be the faces of $\mathcal{F}^{-}(a)$ and $\mathcal{F}^{+}(b)$ bounding quadrants with common boundary $r_{a}$, $r_{b}$ (respectively).
Suppose there exists $g\in G$ such that $g(f_{b})\cap f_{a}\neq\emptyset$. Then for any neighborhood $U$ of $x$, there exists a point $y\in\mathcal{R}_{G}\cap U$ totally linked with $a$ and $b$. In particular, $x\in\overline{\mathrm{Fix}}_{G}$.
Note that in the case where $a,b$ are on nonsingular leaves, the “faces” are simply the leaves through $a$ and $b$, and so the assumptions are satisfied whenever $a$ and $b$ are in the same Smale class.
@@
Let $\alpha$ and $\beta$ be elements fixing $a$ and $b$, respectively. Up to replacing with powers and inverses, we assume they fix all rays of ±$\mathcal{F}^{\pm}(a)$ and ±$\mathcal{F}^{\pm}(b)$ (respectively) and are expanding on $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{+}(b)$. Let $I$ be a small interval of leaves of $\mathcal{L}^{+}$ containing, in its interior, the face $f_{a}$ of $\mathcal{F}^{+}(a)$ such that $g(f_{b})\cap f_{a}\neq\emptyset$ as in the hypotheses of the Proposition. Choose this interval small enough so that $I$ lies inside the saturation of $U$ by $\mathcal{F}^{+}$. Then, for any large enough $n,m>0$, we will have that $g\beta^{n}(I)$ consists of the leaves passing through a small interval of $\mathcal{F}^{-}(a)$, or of $f_{a}$ in the case where $a$ is singular. Thus $\alpha^{m}g\beta^{n}(I)\subset I$. We conclude that $h=\alpha^{m}g\beta^{n}$ fixes a leaf of $\mathcal{F}^{-}$ in $I$.
@@
Suppose $a,b\in\mathrm{Fix}_{G}$ are totally linked. Let $z_{1}$ and $z_{2}$ be the intersections of ±$\mathcal{F}^{\pm}(a)\cap\mathcal{F}^{\mp}(b)$, and let $U_{i}$ be a neighborhood of $z_{i}$ in the quadrilateral bounded by ±$\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\mp}(b)$. Then there exists $g\in G$ such that $g(U_{1})\cap U_{2}\cap\mathrm{Fix}_{G}\neq\emptyset$, and preserving the local orientation of the foliations in the rectangle bounded by ±$\mathcal{F}^{\pm}(a)$ and ±$\mathcal{F}^{\pm}(b)$.

### v2

@@ 2.1. Structure of the bifoliated planes
Here is where a little extra care is needed compared to the case of transitive Anosov-like actions. The assertion above is true for any $l^{\prime\prime}$ in $D(l)$ fixed by some non trivial element $h\in G$. So, thanks to Axiom (A2), we can find such a leaf $l^{\prime\prime}$ that intersects $\mathcal{F}^{+}(x)$. Calling $x^{\prime\prime}$ the fixed point of $h$ on $l^{\prime\prime}$, we will first show that $\mathcal{F}^{+}(x^{\prime\prime})$ intersects $l^{\prime}$ [^4] This is the only argument that needs to be added compared to [Bar95a] or [BFM25b], since in the transitive case, one can choose $x^{\prime\prime}$ close to $x$ so that property is automatically satisfied.. Consider the set $S$ of $\mathcal{F}^{+}$ leaves intersecting both $L^{\prime\prime}$ and $l^{\prime}$. If $\mathcal{F}^{+}(x^{\prime\prime})$ is not in this set, then it means that there is a leaf $l_{1}^{+}$ in $\partial S$ that separates $\mathcal{F}^{+}(x^{\prime\prime})$ from $S$. Since $\alpha(l^{\prime})$ and $\omega(l^{\prime})$ are not in $S$ either, there are also leaves $l_{\alpha}^{+},l_{\omega}^{+}\subset\partial S$ that separates $S$ from $\alpha(l^{\prime})$ and $\omega(l^{\prime})$ respectively (we may a priori have $\alpha(l^{\prime})\subset\partial S$ in which case $\alpha(l^{\prime})=l_{\alpha}^{+}$ and same for $\omega(l^{\prime})$). Then $l_{1}^{+}$ is not separated with either $l_{\alpha}^{+}$ or $l_{\omega}^{+}$, contradicting the fact that $\mathcal{L}^{+}$ is homeomorphic to $\mathbb{R}$. Hence, $\mathcal{F}^{+}(x^{\prime\prime})$ intersects $l^{\prime}$ and one finishes the proof as in [Bar95a] or [BFM25b]. ∎
@@
This was proved for flows by Barbot [Bar95a, Théorème 2.7], and for transitive Anosov-like actions in [BFM25b, Proposition 2.7], the proof there does not use the full strength of transitivity and carries over in our setting without modification. In the case of flows, this is equivalent to the fact that any (pseudo)-Anosov flow with trivial orbit space is the suspension flow of a hyperbolic linear map of the torus.
@@ 2.2. Axiom (A4) and non-separated leaves
Axiom (A4) states that any non-separated leaf is fixed by some nontrivial element of $G$. Our next lemma improves this to show that for any maximal set of pairwise non-separated leaves, there exists a single (nontrivial) element $g\in G$ fixing all the leaves in that set. In particular this shows that Axiom (A4) implies the (a priori stronger) corresponding axiom in [BFM25b].
@@
Take $l_{0}\in S$ and $g\in G$ fixing $l_{0}$. If $g$ is trivial there is nothing to show; so we assume $g$ nontrivial. By Axiom (A1), there exists a unique $x_{0}\in l_{0}$ fixed by $g$. Up to taking a power, we further assume that that $g$ fixes all the rays of $\mathcal{F}^{\pm}(x_{0})\smallsetminus\{x_{0}\}$.
@@
Pick $l_{0}^{\prime}\in S$ another leaf. By Observation 2.13, there exists $f_{0}\in\mathcal{F}^{-}$ that makes a perfect fit with $l_{0}$ and separates $l_{0}$ from $l^{\prime}_{0}$. As $g$ fixes every ray of $x_{0}$, it also fixes $f_{0}$ and Axiom (A1) again gives the existence of a fixed point $y_{0}\in f_{0}$. Let $r_{0}$ be the ray of $\mathcal{F}^{+}(y_{0})\smallsetminus\{y_{0}\}$ that is contained in the side of $f_{0}$ that contains $l^{\prime}_{0}$. Call $S_{0}$ the subset of $S$ on that same side of $f_{0}$. So in particular, $l^{\prime}_{0}\in S_{0}$.
@@
Let $z$ be as above. There exists a unique leaf $f_{1}\in\mathcal{F}^{-}$ such that it, or one of its face in the singular case, is in $\partial(P\smallsetminus\mathcal{F}^{-}(r_{0}))$ and separates $z$ from $x_{0}$. We claim that $f_{1}\cap r_{0}=\emptyset$. This is automatic if $f_{1}$ is non-singular, but assume not and call $t_{0}=f_{1}\cap r_{0}$. Then, by construction, for all $t>t_{0}$ in $r_{0}$ (where >> designate the positive direction of the ray), we have $\mathcal{F}^{-}(t)\cap S_{0}=\emptyset$. However, since $S$ consists of pairwise non-separated leaves, there exists $l^{+}$ that intersects both $f_{0}$ and $\mathcal{F}^{-}(z)$. But (up to replacing $g$ by $g^{-1}$) the sequence $g^{n}l^{+}$ contains $\mathcal{F}^{+}(y_{0})$ in its limit as $n\to+\infty$. In particular, we would get $t>t_{0}$ in $r_{0}$ such that $\mathcal{F}^{-}(t)\cap S_{0}\neq\emptyset$, a contradiction.
@@ 2.3. Product and scalloped regions
Suppose $U$ is an infinite product region in $P$, bounded by an interval $I$. Up to switching the labels ++ and -, we assume $I$ lies in a leaf of $\mathcal{F}^{+}$. By axiom (A2) there is a leaf $l^{-}$ of $\mathcal{F}^{-}$ intersecting $U$ and fixed by some $g\in G$. Let $x$ be the unique fixed point of $g$ on $l^{-}$. Then $\bigcup_{k\in\mathbb{Z}}g^{k}(U)$ contains an infinite product region bounded by $\mathcal{F}^{+}(x)$. Let $y_{i}$ be points along $\mathcal{F}^{+}(x)$ such that the sequence $y_{i}$ eventually leaves every compact set, and for each $i$ let $r_{i}$ be a ray of $\mathcal{F}^{-}(y_{i})$ contained in $U$.
@@
Lemma 2.19 (Lemma 2.23 and Prop. 2.24 in [BFM25b], see also [Fen95]).
@@
The proof of Lemma 2.23 and Prop. 2.24 in [BFM25b] applies directly in this setting, given Proposition 2.18 and Proposition 2.12. ∎
We introduce another definition (from [BFM25b]).
@@
It turns out that infinite lines of lozenges are special: They are part of what is called a scalloped region. A scalloped region is a trivially foliated open set $U\subset P$ whose boundary $\partial U$ consists in four infinite families of pairwise non-separated leaves. In particular, a scalloped region can be realized as two distinct bi-infinite lines of lozenges: one sharing sides along leaves of $\mathcal{F}^{+}$ and the other sharing $\mathcal{F}^{-}$-sides. See Figure 2. A formal definition can be found in e.g., [BFM25b, Definition 2.31].
@@
The fact that infinite lines are contained in bi-infinite lines is given by Lemma 2.32 of [BFM25b], the second assertion is given by Lemma 2.33 of [BFM25b]. Note that neither of these proofs use the transitivity axioms, hence apply directly in our more general setting. ∎
@@ 2.4. Finite branching
[…]
@@
By definition the set of singular points is discrete and it is also closed since any nonsingular point has a product foliated neighborhood. Thus, we need only treat the case of pivots.
Suppose for contradiction that $x_{n}$ is a sequence of pivot points converging to some point $x$. After passing to a subsequence, we can assume that $x_{n}$ lie in a single quadrant $Q$ of $x$, and passing to the tail end we assume that all are in a trivially foliated subset of this quadrant, in particular they are nonsingular. Up to taking a further subsequence (and potentially relabelling the foliations) we may assume that the leaves $\mathcal{F}^{+}(x_{n})$ are always the shared sides of two lozenges.
For concreteness, fix a local orientation on $Q$ in a product-foliated neighborhood of $x$ so that $Q$ is the upper-left quadrant of $x$ and the local foliations are represented with $\mathcal{F}^{+}$ as vertical and $\mathcal{F}^{-}$ horizontal. After passing to a further subsequence, we can assume that the pivots all have the same orientation, that is to say that the perfect fits of all the $\mathcal{F}^{+}(x_{n})$ leaves are either all above $\mathcal{F}^{-}(x)$ or all below $\mathcal{F}^{-}(x)$. We will show that each case contradicts Lemma 2.7
Suppose as a first case that the perfect fits are above. Note that the fact that all $x_{i}$ are in a trivially foliated neighborhood of $x$ in $Q$ implies that each of $\mathcal{F}^{\pm}(x_{i})$ intersects $\mathcal{F}^{\mp}(x)$ (respectively). Take $n$ large enough so that $x_{n}$ lies in the rectangle bounded by $\mathcal{F}^{\pm}(x_{1})$ and $\mathcal{F}^{\pm}(x)$. Then the perfect fit formed by $\mathcal{F}^{+}(x_{1})$ and the side of the rightmost lozenge with corner $x_{1}$ is in the top left quadrant of $x_{n}$, so, by [BFM25b, Lemma 2.29] $x_{n}$ cannot have a lozenge there, a contradiction. Similarly, if the perfect fits are all below, $x_{n}$ cannot have a lozenge in its lower left quadrant, again a contradiction. ∎
@@ 3. Smale classes and topological transitivity
In this section, we generalize the Smale order and basic sets for Anosov flows to the context of an Anosov-like action of a group $G$ on bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$. In particular, we show how the classical Smale order and basic sets (see [Sma67]) can be explicitly detected in the orbit space, that Smale classes have a product structure, and we prove Theorems 1.5 and 1.7. Later, in Section 4, we will also recover the analogue of tori separating basic sets from the orbit space.
@@
The regular set $\mathcal{R}_{G}$ is the subset of $\overline{\mathrm{Fix}}_{G}$ consisting of points $x\in\overline{\mathrm{Fix}}_{G}$ such that no leaf of $\mathcal{F}^{\pm}(x)$ is a singular leaf. Equivalently, $\mathcal{R}_{G}$ is obtained from $\overline{\mathrm{Fix}}_{G}$ by removing all prongs and their leaves.
@@
However, even for actions coming from orbit spaces of pseudo-Anosov flows on compact manifolds, there may be differences between chain-recurrent and non-wandering sets. In this case $\overline{\mathrm{Fix}}_{G}$ corresponds to the projection of the non-wandering set. The non-wandering set can be strictly smaller than the chain recurrent set, as shown by the examples given in Propositions 8.8 and 8.9.
For Anosov-like actions, the interesting dynamics happens in the set $\overline{\mathrm{Fix}}_{G}$. Working in the regular set $\mathcal{R}_{G}$ is necessary to deal with the unusual behavior that arise in the presence of prongs, but it may happen that the closure of $\mathcal{R}_{G}$ misses parts (or even all, see Proposition 8.10) of $\overline{\mathrm{Fix}}_{G}$, and therefore misses parts of the dynamics. Thus, we introduce some terminology for points of $\overline{\mathrm{Fix}}_{G}\smallsetminus\bar{\mathcal{R}}_{G}$ and we will keep track of them separately.
Definition 3.3.
A point $p\in P$ is called an isolated prong singularity if $p$ is a singularity of the foliations and an isolated point of $\overline{\mathrm{Fix}}_{G}$. Equivalently, isolated prong singularities are the points in $\overline{\mathrm{Fix}}_{G}\smallsetminus\overline{\mathcal{R}}_{G}$.
@@
We extend this relation to isolated prong singularities, by saying $x\sim_{G}y$ if $x=gy$ for some $g\in G$.
@@
In the case of the orbit space of an Anosov flow, the relation $\leq_{G}$ is defined and a transitive relation on all of $\overline{\mathrm{Fix}}_{G}$, giving the classical Smale order, and the Smale classes are basic sets ([Sma67], or see e.g., [FH19, Section 5.3]). Transitivity of the relation $\leq_{G}$ on $\overline{\mathrm{Fix}}_{G}$ also holds whenever $\mathcal{F}^{\pm}$ are true foliations, without singularities, by Lemma 3.5 below. However, in the presence of prongs, the property of having nonempty intersection is not open in the space of pairs of leaves in $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$ and transitivity may fail for points in $\overline{\mathrm{Fix}}_{G}$ – see Proposition 8.7. The regular set $\mathcal{R}_{G}$ is the set where $\leq_{G}$ always induces a partial order on Smale classes, as we show in the next lemma.
@@
The relation $\leq_{G}$ is reflexive and transitive on $\mathcal{R}_{G}$. Consequently, the relation $\sim_{G}$ is an equivalence relation, and $\leq_{G}$ induces a partial order on the Smale classes that do not consist of the orbit of an isolated prong singularity.
@@
In order to more succinctly differentiate between Smale classes whose elements are in $\mathcal{R}_{G}$ from the ones containing only isolated prong singularities, we introduce the following terminology:
Definition 3.6.
A Smale class $\Lambda$ is called regular if $\Lambda\subset\mathcal{R}_{G}$. It is called singular if $\Lambda$ is the orbit of an isolated prong singularity.
Since we have an order on regular Smale classes, we introduce the following terminology, which will be important in Section 6.
Definition 3.7 (Extremal classes).
We call a Smale class extremal if it is regular and either minimal or maximal for the Smale order.
Our next goal is to give an equivalent characterization of Smale classes that is easier to work with. For this, we use the following definition from [BFM25b]
@@
[…]
As in the classical case of hyperbolic basic sets, we will now show that the Smale classes of an Anosov-like action have a product structure, in the same sense as hyperbolic sets – see e.g., [FH19, Definition 6.2.6].
@@
Let $a,b\in\overline{\mathrm{Fix}}_{G}$ satisfy $x=\mathcal{F}^{+}(a)\cap\mathcal{F}^{-}(b)\neq\emptyset$. Assume that $x$ is not singular. Let $r_{a}$ and $r_{b}$ be the rays of $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{-}(b)$ (respectively) containing $x$, and let $f_{a},f_{b}$ be the faces of $\mathcal{F}^{-}(a)$ and $\mathcal{F}^{+}(b)$ bounding quadrants with common boundary $r_{a}$, $r_{b}$ (respectively).
If there exists $g\in G$ such that $g(f_{b})\cap f_{a}\neq\emptyset$, then for any neighborhood $U$ of $x$, there exists a point $y\in\mathcal{R}_{G}\cap U$ totally linked with $a$ and $b$. In particular, $x\in\overline{\mathrm{Fix}}_{G}$. [^6] If $x$ was singular, then we automatically would have that $x\in\overline{\mathrm{Fix}}_{G}$ and is totally linked with both $a$ and $b$, but may not have any regular fixed points totally linked with both $a$ and $b$.
Note that in the case where $a,b$ are on nonsingular leaves, the “faces” are simply the leaves through $a$ and $b$, and so the assumptions are satisfied whenever $a$ and $b$ are in the same Smale class, giving the statement of Proposition 3.9.
@@
Let $\alpha$ and $\beta$ be elements fixing $a$ and $b$, respectively. Up to replacing with powers and inverses, we assume they fix all rays of $\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\pm}(b)$ (respectively) and are expanding on $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{+}(b)$. Let $I$ be a small interval of leaves of $\mathcal{L}^{+}$ containing, in its interior, the face $f_{a}$ of $\mathcal{F}^{+}(a)$ such that $g(f_{b})\cap f_{a}\neq\emptyset$ as in the hypotheses of the Proposition. Choose this interval small enough so that $I$ lies inside the saturation of $U$ by $\mathcal{F}^{+}$. Then, for any large enough $n,m>0$, we will have that $g\beta^{n}(I)$ consists of the leaves passing through a small interval of $\mathcal{F}^{-}(a)$, or of $f_{a}$ in the case where $a$ is singular. Thus $\alpha^{m}g\beta^{n}(I)\subset I$. We conclude that $h=\alpha^{m}g\beta^{n}$ fixes a leaf of $\mathcal{F}^{-}$ in $I$.
@@
Suppose $a,b\in\mathrm{Fix}_{G}$ are totally linked. Let $z_{1}$ and $z_{2}$ be the intersections of $\mathcal{F}^{\pm}(a)\cap\mathcal{F}^{\mp}(b)$, and let $U_{i}$ be a neighborhood of $z_{i}$ in the quadrilateral bounded by $\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\mp}(b)$. Then there exists $g\in G$ such that $g(U_{1})\cap U_{2}\cap\mathrm{Fix}_{G}\neq\emptyset$, and preserving the local orientation of the foliations in the rectangle bounded by $\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\pm}(b)$.